# Revisiting Cross-Diffusion for Overcrowding Dispersal in Interacting Species System

### v1

@@ 5 Connection with $H^{-1}$-like theory in the homogeneous case
$\pi=0\hbox{ on }\Gamma_{N},\quad g=0\hbox{ on }\Gamma_{D}\quad\hbox{ and }\quad V\equiv 0.$
@@
$[f,g]_{\Omega}=\langle f_{1},g_{1}\rangle_{1,\Omega}+\langle f_{2},g_{2}\rangle_{k,\Omega},\quad\hbox{ for any }f,g\in[H^{-1}_{\Gamma_{D}}(\Omega)]^{2}.$
Our main result in this section concerns the characterization of $\partial E$ in terms of cross-diffusion system like in (3.37).
@@
For any $(\rho,f)\in[H^{-1}_{\Gamma_{D}}(\Omega)]^{2}\times[H^{-1}_{\Gamma_{D}}(\Omega)]^{2}.$ Then, $f\in\partial E(\rho)$ if and only if $\rho\in[L^{2}_{+}(\Omega)]^{2}$ and $\rho$ solves the following system
(5.69) $\left\{\begin{array}[]{ll}\left.\begin{array}[]{ll}-\nabla\cdot(\sigma_{k}\>\nabla\eta_{k})=f_{0k}+\nabla\cdot\overline{f}_{k},\quad\hbox{ for }k=1,2\\ \\ \rho_{1},\>\rho_{2}\geq 0,\>\frac{\eta_{1}}{\alpha_{1}}\vee\frac{\eta_{2}}{\alpha_{2}}\in\partial\beta(x,\alpha_{1}\rho_{1}+\alpha_{2}\rho_{2})\\ \\ \rho_{1}\left(\frac{\eta_{2}}{\alpha_{2}}-\frac{\eta_{1}}{\alpha_{1}}\right)^{+}=\rho_{2}\left(\frac{\eta_{1}}{\alpha_{1}}-\frac{\eta_{2}}{\alpha_{2}}\right)^{+}=0\\ \end{array}\right\}&\hbox{ in }\Omega,\\ \\ (\sigma_{k}\>\nabla\eta_{k}+\overline{f}_{k})\cdot\nu=0&\hbox{ on }\Gamma_{N_{k}}\\ \\ \eta_{k}=g_{k}&\hbox{ on }\Gamma_{D_{k}},\end{array}\right.$
in the sense that, there exists $\eta_{k}\in H^{1}(\Omega),$ such that $\eta_{k}=g_{k}$ on $\Gamma_{D_{k}},$ for $k=1,2,$ and we have
  • $\rho_{1}\left(\frac{\eta_{2}}{\alpha_{2}}-\frac{\eta_{1}}{\alpha_{1}}\right)^{+}=\rho_{2}\left(\frac{\eta_{1}}{\alpha_{1}}-\frac{\eta_{2}}{\alpha_{2}}\right)^{+}=0,$ a.e. in $\Omega,$
  • $\frac{\eta_{1}}{\alpha_{1}}\vee\frac{\eta_{2}}{\alpha_{2}}\in\partial\beta(x,\alpha_{1}\rho_{1}+\alpha_{2}\rho_{2})$, a.e. in $\Omega,$
  • For each $k=1,2,$
    $\int_{\Omega}(\sigma_{k}\>\nabla\eta_{k}+\chi_{k})\cdot\nabla\xi_{k}=\int_{\Omega}\mu_{k}\>\xi_{k}-\langle\pi_{k},\xi_{k}\rangle_{\Gamma_{k}}.$
    for any $\xi_{k}\in H^{1}_{\Gamma_{{}_{k}}}(\Omega).$
@@
where $\mathcal{A}$ is the operator define in $[H^{-1}_{\Gamma_{D}}(\Omega)]^{2},$ by $f\in\mathcal{A}(\rho)$ if and only if $f\in[H^{-1}_{\Gamma_{D}}(\Omega)]^{2},$ $\rho\in[L^{2}(\Omega)_{+}]^{2}$ and $\rho$ is a solution of (5.69), in the sense of Proposition 5. Thanks to Corollary 1, for any $f\in[H^{-1}_{\Gamma_{D}}(\Omega)]^{2},$ there exists $\rho\in[L^{2}_{+}(\Omega)]^{2},$ such that $f=\rho+\mathcal{A}(\rho).$ On the other hand, since $\mathcal{E}$ is convex and l.s.c. in $[H^{-1}_{\Gamma_{D}}(\Omega)]^{2},$ the operator $\partial E$ defines a maximal monotone graph in $[H^{-1}_{\Gamma_{D}}(\Omega)]^{2}.$ So, for the proof of (5.70), it is enough to prove that $\mathcal{A}\subseteq\partial\mathcal{E}.$ To this aim, we consider $f\in[H^{-1}_{\Gamma_{D}}(\Omega)]^{2}$ and $\rho\in[L^{2}(\Omega)_{+}]^{2}$ be such that $f\in\mathcal{A}(\rho),$ and we prove that
(5.71) $[f,z-\rho]_{\Omega}\leq\int_{\Omega}\beta_{\alpha}[z]-\int_{\Omega}\beta_{\alpha}[\rho]+,\quad\hbox{ for any }z\in[L^{2}(\Omega)_{+}]^{2}.$
@@
$\phi^{z}_{k}=\hbox{argmin}_{\omega\in L^{2}(\Omega)^{N}}\left\{\int_{\Omega}|\sigma_{k}\>\omega|^{2}\>:\>-\nabla\cdot(\sigma_{k}\>\omega)=z_{k}\hbox{ in }\Omega\hbox{ and }(\sigma_{k}\>\omega)\cdot\nu=0\hbox{ on }\Gamma_{N_{k}}\right\},$
@@
$\phi^{\rho}_{k}=\hbox{argmin}_{\omega\in L^{2}(\Omega)^{N}}\left\{\int_{\Omega}|\sigma_{k}\>\omega|^{2}\>:\>-\nabla\cdot(\sigma_{k}\>\omega)=\rho_{k}\hbox{ in }\Omega\hbox{ and }(\sigma_{k}\>\omega)\cdot\nu=0\hbox{ on }\Gamma_{N_{k}}\right\}.$
@@
$\displaystyle$== $\displaystyle\int_{\Omega}\sigma_{1}\>\nabla\eta_{1}\cdot(\phi^{z}_{1}+\phi^{\rho}_{1})+\int_{\Omega}\sigma_{2}\>\nabla\eta_{2}\cdot(\phi^{z}_{2}+\phi^{\rho}_{2})$
$\displaystyle$== $\displaystyle\int_{\Omega}(\eta_{1}\>z_{1}+\eta_{2}\>z_{2})-\int_{\Omega}(\eta_{1}\>\rho_{1}+\eta_{2}\>\rho_{2})$
Moreover, since $\frac{\eta_{1}}{\alpha_{1}}\vee\frac{\eta_{2}}{\alpha_{2}}\in\partial\beta(x,\alpha_{1}\>\rho_{1}+\alpha_{2}\>\rho_{2}),$ a.e. in $\Omega,$ we obtain
$\displaystyle[f,z-\rho]_{\Omega}$ $\displaystyle$== $\displaystyle\int_{\Omega}(\eta_{1}\>z_{1}+\eta_{2}\>z_{2})-\int_{\Omega}(\alpha_{1}\>\rho_{1}+\alpha_{2}\>\rho_{2})\frac{\eta_{1}}{\alpha_{1}}\vee\frac{\eta_{2}}{\alpha_{2}}$
@@
: Since $\partial\mathcal{E}$ is a maximal monotone graph, the existence of a strong solution of (2.21) follows directly from the well-established theory of nonlinear semigroups in Hilbert spaces (see [9] for details). This allows to conclude that Theorem 2 holds true as a direct consequence of Proposition 5. $\square$

### v2

@@ 5 Connection with $H^{-1}$-like theory in the homogeneous case
$\pi=0\hbox{ on }\Gamma_{N}\quad\hbox{ and }\quad g=0\hbox{ on }\Gamma_{D}.$
@@
$[f,g]_{\Omega}=\langle f_{1},g_{1}\rangle_{1,\Omega}+\langle f_{2},g_{2}\rangle_{2,\Omega},\quad\hbox{ for any }f,g\in[H^{-1}_{\Gamma_{D}}(\Omega)]^{2}.$
Our main result in this section concerns the characterization of $\partial E$ in terms of cross-diffusion system like in (3.51). As we’ll see this operator is closely connected to the stationary problem
(5.86) $\left\{\begin{array}[]{ll}\left.\begin{array}[]{ll}-\nabla\cdot(\sigma_{k}\>\nabla\eta_{k}+\overline{f}_{k})=f_{0k},\quad\hbox{ for }k=1,2\\ \\ \eta_{k}\in\partial\beta(.,\rho_{1}+\rho_{2})+\partial I\!\!I_{[0,\infty)}(\rho_{k})\\ \end{array}\right\}&\hbox{ in }\Omega,\\ \\ (\sigma_{k}\>\nabla\eta_{k}+\overline{f}_{k})\cdot\nu=0&\hbox{ on }\Gamma_{N_{k}}\\ \\ \eta_{k}=0&\hbox{ on }\Gamma_{D_{k}},\end{array}\right.\quad k=1,2,$
for a given $f=(f_{1},f_{2})\in H^{-1}_{\Gamma_{D_{1}}}(\Omega)\times H^{-1}_{\Gamma_{D_{2}}}(\Omega),$ and for each $k=1,2,$ $(f_{0k},\overline{f}_{k})\in L^{2}(\Omega)\times L^{2}(\Omega)^{N}$ is the couple associated with $f_{k}\in H^{-1}_{\Gamma_{D_{k}}}(\Omega),$ $f_{k}=f_{0k}+\nabla\cdot\overline{f}_{k}.$
@@
For any $(\rho,f)\in[H^{-1}_{\Gamma_{D}}(\Omega)]^{2}\times[H^{-1}_{\Gamma_{D}}(\Omega)]^{2},$ $f\in\partial E(\rho)$ if and only if $\rho:=(\rho_{1},\rho_{2})\in[L^{2}_{+}(\Omega)]^{2},$ and for each $k=1,2,$ there exists $\eta_{k}\in H^{1}_{\Gamma_{D_{k}}}(\Omega),$ such that
$\eta_{1}\vee\eta_{2}=:\tilde{\eta}\in\partial\beta(x,\rho_{1}+\rho_{2}),\>\quad\eta_{k}-\tilde{\eta}\in\partial I\!\!I_{[0,\infty)}(\rho_{k}),\quad\hbox{ a.e. in }\Omega,$
and
$\int_{\Omega}\sigma_{k}\>\nabla\eta_{k}\cdot\nabla\xi_{k}=\langle f_{k},\xi_{k}\rangle_{k,\Omega}-\langle\pi_{k},\xi_{k}\rangle_{\Gamma_{k}},\quad\hbox{ for any }\xi_{k}\in H^{1}_{\Gamma_{{}_{k}}}(\Omega).$
@@
where $\mathcal{A}$ is the operator define in $[H^{-1}_{\Gamma_{D}}(\Omega)]^{2},$ by $f\in\mathcal{A}(\rho)$ if and only if $f\in[H^{-1}_{\Gamma_{D}}(\Omega)]^{2},$ $\rho\in[L^{2}(\Omega)_{+}]^{2}$ and $\rho$ is a solution of (5.86), in the sense of Proposition 5.
Thanks to classical theory of maximal monotone graphs (cf. [10]), since $\mathcal{E}$ is convex and l.s.c. in $[H^{-1}_{\Gamma_{D}}(\Omega)]^{2},$ the operator $\partial E$ defines a maximal monotone graph in $[H^{-1}_{\Gamma_{D}}(\Omega)]^{2}.$ Since E is both convex and lower semi-continuous in $[H^{-1}_{\Gamma_{D}}(\Omega)]^{2},$ $\partial E$ itself defines a maximal monotone graph within the same space. Furthermore, Corollary 3 guarantees that for any $f\in[H^{-1}_{\Gamma_{D}}(\Omega)]^{2},$ there exists $\rho\in[L^{2}_{+}(\Omega)]^{2},$ such that $f=\rho+\mathcal{A}(\rho).$ To establish the identification in equation (5.87), it suffices to prove that $\mathcal{A}\subset\partial\mathcal{E}$. Indeed, proving this inclusion, implies that $\mathcal{A}$ is a maximal monotone graph contained within $\partial\mathcal{E}.$ Consequently, the two graphs must coincide. To this aim, we consider $f\in[H^{-1}_{\Gamma_{D}}(\Omega)]^{2}$ and $\rho\in[L^{2}(\Omega)_{+}]^{2}$ be such that $f\in\mathcal{A}(\rho),$ and we prove that
(5.88) $[f,z-\rho]_{\Omega}\leq\int_{\Omega}\beta[z]-\int_{\Omega}\beta[\rho],\quad\hbox{ for any }z\in[L^{2}(\Omega)_{+}]^{2}.$
@@
$\phi^{z}_{k}=\hbox{argmin}_{\omega\in L^{2}(\Omega)^{N}}\left\{\int_{\Omega}\sigma_{k}\>|\omega|^{2}\>:\>-\nabla\cdot(\sigma_{k}\>\omega)=z_{k}\hbox{ in }\Omega\hbox{ and }\sigma_{k}\>\omega\cdot\nu=0\hbox{ on }\Gamma_{N_{k}}\right\},$
@@
$\phi^{\rho}_{k}=\hbox{argmin}_{\omega\in L^{2}(\Omega)^{N}}\left\{\int_{\Omega}\sigma_{k}\>|\omega|^{2}\>:\>-\nabla\cdot(\sigma_{k}\>\omega)=\rho_{k}\hbox{ in }\Omega\hbox{ and }\sigma_{k}\>\omega\cdot\nu=0\hbox{ on }\Gamma_{N_{k}}\right\}.$
@@
$\displaystyle$== $\displaystyle\int_{\Omega}\sigma_{1}\>\nabla\eta_{1}\cdot(\phi^{z}_{1}-\phi^{\rho}_{1})+\int_{\Omega}\sigma_{2}\>\nabla\eta_{2}\cdot(\phi^{z}_{2}-\phi^{\rho}_{2})$
$\displaystyle$== $\displaystyle\int_{\Omega}\eta_{1}\>(z_{1}-\rho_{1})+\int_{\Omega}\eta_{2}\>(z_{2}-\rho_{2})$
$\displaystyle$== $\displaystyle\int_{\Omega}(\eta_{1}\>z_{1}+\eta_{2}\>z_{2})-\int_{\Omega}(\eta_{1}\>\rho_{1}+\eta_{2}\>\rho_{2}).$
Remember that $(\eta_{1},\eta_{2})\in\partial\beta(.,(\rho_{1},\rho_{2})),$ a.e. in $Q.$ So $\eta_{1}\rho_{1}+\rho_{2}\eta_{2}=(\rho_{1}+\rho_{2})\eta_{1}\vee\eta_{2}$ and ${\eta_{1}}\vee\eta_{2}\in\partial\beta(x,\rho_{1}+\rho_{2}),$ a.e. in $\Omega$. This implies that
$\displaystyle[f,z-\rho]_{\Omega}$ $\displaystyle$== $\displaystyle\int_{\Omega}(\eta_{1}\>z_{1}+\eta_{2}\>z_{2})-\int_{\Omega}(\rho_{1}+\rho_{2}){\eta_{1}}\vee\eta_{2}$
@@
: The proof follows directly from Proposition 5 and the fact that $h=f-[\nabla\cdot(\rho\>V)]\in L^{2}(0,T;[H^{-1}_{\Gamma_{D}}(\Omega)]^{2}).$ Then, the proof is a direct consequence of the definition of weak solution of (2.15) and the characterization of $\partial\mathcal{E}$ in terms of PDE in Proposition 5. $\square$